# Chinese Academy of Sciences, Beijing 100190, China

### v2

@@ 4.2 Results of BBN Abundances
where the subscript $\xi_{\nu}=0$ indicates results obtained in the SM case using actual density matrices, as detailed in Table 1. These fitting functions facilitate further studies, yielding fractional errors smaller than $0.01\%$ within the range $-0.1\leq\xi_{\nu}\leq 0.1$.  [^3] The coefficients of our fitting functions differ slightly from those in Pitrou:2018cgg ; Froustey:2024mgf , which is mainly due to the different neutron lifetimes we have chosen. In addition, our fitting function is more precise.
@@ 5.1.1 Effects on Neutrino Abundance
$\overset{\delta\varrho=0}{\simeq}\frac{\sum\limits_{i}m_{i}}{93.12\,{\rm eV}}\frac{{\rm Li}_{3}(-\exp(-\xi_{\nu}))+{\rm Li}_{3}(-\exp(\xi_{\nu}))}{3\zeta(3)/2},$
@@
The present neutrino abundance is further modified by spectral distortions of neutrinos and antineutrinos, though the integrals over their distributions cannot be carried out analytically in this case. The effect of primordial neutrino asymmetries becomes clearer when spectral distortions are neglected, allowing an analytical evaluation of the integrals, as shown in the second row of (90). The first factor, $\frac{\sum\limits_{i}m_{i}}{93.12\,{\rm eV}}$, represents the usual neutrino relic abundance in the SM, while the second factor accounts for the effects of neutrino asymmetries, with ${\rm Li}_{j}(x)$ being the polylogarithms of order $j$ and $\zeta(x)$ the Riemann zeta function. In the SM case, ${\rm Li}_{3}(-1)=3\zeta(3)/4$, so the second factor equals 1. For non-zero $\xi_{\nu}$, the second factor depends on the absolute value of $\xi_{\nu}$, with larger $\xi_{\nu}$ leading to a larger factor and, consequently, a larger neutrino abundance. However, the corrections from primordial neutrino asymmetries are small. For instance, the relative difference in $\Omega_{\nu}h^{2}$ for $\xi_{\nu}=0.2$ is only $1.54\%$. Similarly, the corrections to the neutrino abundance from spectral distortions of neutrinos and antineutrinos are negligible.
@@ 5.1.2 Effects on the Neutrino Boltzmann Hierarchies
[…]
@@
[…]
@@ 5.1.3 Effects on the BAO
$\dot{\delta}_{\gamma}$ $=$ $-\frac{4}{3}\theta_{\gamma}-\frac{2}{3}\dot{h},$
@@
[…]
@@
[…]
@@ 6.1 Observational Constraints on Primordial Neutrino Asymmetry
$\mathbf{\Lambda}$CDM + $\mathbf{\xi_{\nu}}$ $0.023\pm 0.012$ $2.200\pm 0.011$ $0.1207\pm 0.0009$
$\mathbf{\Lambda}$CDM + $\mathbf{\xi_{\nu}}$ + $\mathbf{m_{\nu_{1}}}$ $0.024\pm 0.012$ $2.200\pm 0.011$ $0.1205\pm 0.0009$ 0.131
@@
To extract constraints from these datasets, we perform MCMC analyses using the Montepython engine Audren:2012wb ; Brinckmann:2018cvx . For the analysis, $N_{\rm eff}$ is derived from Eq. (63) as discussed in Section 3. The predicted helium and deuterium abundances from BBN are calculated using our fitting formulas in Eq. (86). As shown in Fig. 10, the impact of neutrino spectral distortions on the CMB angular power spectrum and BAO in the matter power spectrum is minor. Therefore, for simplicity, (anti)neutrino spectral distortions are neglected in the MCMC analysis. For the CMB and BAO datasets, the parameter set includes the six independent parameters of the $\Lambda$CDM model, along with the neutrino degeneracy parameter $\xi_{\nu}$ and the mass of the lightest neutrino in the normal hierarchy $m_{\nu_{1}}$, i.e., $\{\omega_{b},\omega_{cdm},\theta_{s},A_{s},n_{s},\tau_{\mathrm{reio}},\xi_{\nu},m_{\nu_{1}}\}$. Here, $\theta_{s}$ is the sound horizon, $A_{s}$ is the amplitude of primordial curvature perturbations at the pivot scale $k=0.05\,\mathrm{Mpc}^{-1}$, $n_{s}$ is the scalar spectral index, and $\tau_{\mathrm{reio}}$ is the optical depth to reionization. For the BBN dataset, we consider only $\{\omega_{b},\xi_{\nu}\}$, as the predicted light element abundances from BBN are largely independent of other cosmological parameters within the relevant range.
@@
Using the specified parameters and datasets, we modified the MontePython sampler and CLASS codes to perform the MCMC analysis. All chains converged with a Gelman-Rubin criterion of $R-1<0.01$. The resulting $1\,\sigma$ cosmological constraints are summarized in Table 3, where the $95\%$ confidence limits on the total neutrino mass are also presented, derived from constraints on the mass of the lightest neutrino for the normal hierarchy, $m_{\nu_{1}}$. For the model $\Lambda$CDM + ${\xi_{\nu}}$ + ${m_{\nu_{1}}}$, we present the 1D and 2D marginalized posterior distributions of key parameters that show significant variations across different datasets in Fig. 11, including the baryon abundance $\omega_{b}$, the neutrino degeneracy parameter $\xi_{\nu}$, the cold dark matter abundance $\omega_{cdm}$, the sound horizon $\theta_{s}$, and the mass of the lightest neutrino for the normal hierarchy $m_{\nu_{1}}$.
@@
We find that constraints on other parameters, such as $\omega_{b}$, $\omega_{cdm}$, $n_{s}$, and $m_{\nu_{1}}$, also vary significantly depending on the dataset used, as shown in Fig. 11. For $\omega_{b}$, the discrepancy is mainly attributed to the choice of nuclear reaction rates in the PRIMAT code during the BBN process, which predicts a deuterium abundance lower than current constraints from PDG, leading to a significantly lower $\omega_{b}$ Pitrou:2018cgg ; Burns:2022hkq ; Escudero:2022okz . While constraints on the primordial neutrino asymmetry $\xi_{\nu}$ do not strongly depend on the choice of nuclear reaction rates, as discussed in Escudero:2022okz , a detailed study using different nuclear reaction rates will be presented in our companion paper. The inclusion of BOSS BAO data significantly modifies the constraints on $\omega_{cdm}$, $n_{s}$, and $m_{\nu_{1}}$ due to its sensitivity to the evolution of the late Universe. In particular, the $95\%$ upper limit on the total neutrino mass tightens from $0.414$ eV (BBN + CMB) to $0.131$ eV (BBN + CMB + BAO), as BAO data help break the CMB degeneracy. Recently, DESI BAO observations have placed even stronger constraints on neutrino masses, $\sum m_{\nu_{i}}\lesssim 0.07\,\mathrm{eV}$ DESI:2024mwx . Relevant studies on primordial neutrino asymmetry and neutrino masses using DESI BAO data will be addressed in our companion paper.
Finally, we briefly discuss the impact of different parameter sets on the resulting constraints, particularly the interplay between primordial neutrino asymmetry and total neutrino mass. Using the BBN + CMB + BAO dataset, we present in Fig. 12 the 1D marginalized distributions for three parameter sets: $\Lambda$CDM + $m_{\nu_{1}}$, $\Lambda$CDM + $\xi_{\nu}$, and $\Lambda$CDM + $m_{\nu_{1}}$ + $\xi_{\nu}$. As shown, while different parameter sets lead to significant modifications in constraints on the six $\Lambda$CDM parameters, especially $\omega_{b}$ and $n_{s}$, the resulting constraints on additional parameters are consistent across different parameter sets. The corresponding constraints from different parameter sets are also summarized in Table 3. We find that the constraints on the total neutrino mass are only slightly tightened when $\xi_{\nu}$ is included as a free parameter, while constraints on $\xi_{\nu}$ remain nearly unchanged when $\sum m_{\nu}$ is a free parameter. Therefore, we conclude that constraints on neutrino masses and primordial neutrino asymmetries can be analyzed independently if one is not concerned with constraints on other $\Lambda$CDM parameters. However, if constraints on other $\Lambda$CDM parameters are of interest, a parameter set including both of neutrino mass and primordial neutrino asymmetry must be considered.
@@ 6.2 Constraints on UV Model Parameters
Following Kawasaki:2022hvx , the primordial lepton asymmetry $\eta_{L}$ is defined as the ratio of the difference between the number densities of neutrinos and antineutrinos to the photon number density:
$\eta_{L}\equiv\sum_{\alpha}\frac{n_{\nu_{\alpha}}-n_{\bar{\nu}_{\alpha}}}{s}\simeq 0.105\,\xi_{\nu},$ (102)
where $s=\frac{2\pi^{2}}{45}g_{\star}T_{\gamma}^{3}$ is the entropy density, and $g_{\star}=10.75$ is the relativistic degree of freedom at neutrino decoupling. In the Q-ball decay model, $\eta_{L}$ can also be expressed as:
$\eta_{L}=\frac{3T_{\mathrm{D}}}{4m_{3/2}},$ (103)
where $m_{3/2}$ is the gravitino mass and $T_{\mathrm{D}}$ is the L-balls’ decay temperature, which is a function of other UV model parameters:
@@
×$\hskip 15.0pt\times\left(\frac{g_{*}}{10.75}\right)^{-1/4}\left(\frac{\beta}{6\times 10^{-4}}\right)^{-5/8}\left(\frac{m_{3/2}}{0.5~{}\mathrm{GeV}}\right)^{5/2}\left(\frac{M_{F}}{5\times 10^{6}~{}\mathrm{GeV}}\right)^{-2}\left(\frac{N_{\ell}}{3}\right)^{1/2}\left(\frac{\zeta}{2.5}\right)^{1/2},$ (104)
@@
[…]
@@
[…]
@@ 7 Conclusion
In this work, we address these ambiguities by numerically solving the neutrino decoupling process and analyzing the resulting (anti)neutrino density matrices in the context of BBN, CMB, and LSS. We rederive the quantum kinetic equations (QKEs) for neutrinos and antineutrinos using the Closed Time Path (CTP) formalism for two-point functions. Assuming complete flavor equilibration before neutrino decoupling, where $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{\nu}$, we numerically solve the QKEs in the early universe using an extended version of the public code FortEPiaNO Gariazzo:2019gyi ; Bennett:2020zkv . Our calculations include the effects of neutrino oscillations, matter interactions, FTQED corrections up to $\mathcal{O}(3)$, and comprehensive neutrino-electron and neutrino-neutrino collision terms. The final $N_{\rm eff}$ is then computed based on the resulting density matrices and can be expressed as $N_{\rm eff}\simeq N_{\rm eff}^{\rm SM}+3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right)+0.0102\xi_{\nu}^{2}$, where $N_{\rm eff}^{\rm SM}=3.0440\pm 0.0002$. The second and third terms represent corrections from nonzero $\xi_{\nu}$ in the instantaneous decoupling limit and non-instantaneous effects, respectively. We also confirm that spectral distortions of neutrinos and antineutrinos with nonzero $\xi_{\nu}$ differ due to distinct Pauli-blocking effects, in agreement with preliminary studies Grohs:2016cuu .
@@
Our work establishes a framework for exploring the implications of physics beyond the Standard Model and $\Lambda$CDM for neutrino decoupling and related cosmological processes, providing corresponding constraints from current and future BBN, CMB, and LSS observations. It paves the way for more precise studies of primordial neutrino asymmetries and other new physics scenarios, such as synchronous neutrino oscillations Froustey:2021azz ; Froustey:2024mgf and non-standard neutrino interactions Du:2021idh ; Du:2023upj , driven by increasingly precise cosmological observations.
@@ Acknowledgements.
We thank J. Froustey and C. Pitrou for helpful communications about their relevant work. This work is supported by the National Science Foundation of China under Grants No. 12347105, No. 12375099 and No. 12047503, and the National Key Research and Development Program of China Grant No. 2020YFC2201501, No. 2021YFA0718304. The authors gratefully acknowledge the use of publicly available codes FortEPiaNO Gariazzo:2019gyi ; Bennett:2020zkv , PRIMAT Pitrou:2018cgg , CLASS Lesgourgues:2011re ; Lesgourgues:2011rh and MontePythonAudren:2012wb ; Brinckmann:2018cvx .
@@ A.1 Complete expressions for the collision terms
Finally, the collision terms for the antineutrinos ¯¯$\bar{\mathcal{I}}\equiv\bar{\mathcal{I}}_{\rm sc}+\bar{\mathcal{I}}_{\rm ann}+\bar{\mathcal{I}}_{\nu\nu}$ can be obtained simply by the transformations $\varrho\leftrightarrow\bar{\varrho}$ and $L\leftrightarrow R$.
@@ A.2 Reduced expressions for the comoving collision terms
~$\tilde{\mathcal{I}}_{\rm sc}$ $=$ $\int{\rm d}y_{2}{\rm d}y_{3}\frac{y_{2}}{E_{2}}$
@@
~$\tilde{\mathcal{I}}_{\rm ann}$ $=$ $\int{\rm d}y_{2}{\rm d}y_{3}\frac{y_{3}}{E_{3}}$
@@
~$\tilde{\mathcal{I}}_{\nu\nu}$ $=$ $\frac{1}{4}\int{\rm d}y_{2}{\rm d}y_{3}\,$
@@
Again, the comoving collision terms for the antineutrinos ¯~¯~$\tilde{\bar{\mathcal{I}}}\equiv\tilde{\bar{\mathcal{I}}}_{\rm sc}+\tilde{\bar{\mathcal{I}}}_{\rm ann}+\tilde{\bar{\mathcal{I}}}_{\nu\nu}$ can be obtained simply by the transformations $\varrho\leftrightarrow\bar{\varrho}$ and $L\leftrightarrow R$.

### v3

@@ 4.2 Results of BBN Abundances
where the subscript $\xi_{\nu}=0$ indicates results obtained in the SM case using actual density matrices, as detailed in Table 2. These fitting functions facilitate further studies, yielding fractional errors smaller than $0.01\%$ within the range $0.1\leq\xi_{\nu}\leq 0.1$.  [^3] The coefficients of our fitting functions differ slightly from those in Pitrou:2018cgg ; Froustey:2024mgf , which is mainly due to the different neutron lifetimes we have chosen. In addition, our fitting function is more precise.
@@ 5.1.1 Effects on Neutrino Abundance
$\overset{\delta\varrho=0}{\simeq}\frac{\sum\limits_{i}m_{i}}{93.12\,{\rm eV}}\left(-\frac{{\rm Li}_{3}(-\exp(-\xi_{\nu}))+{\rm Li}_{3}(-\exp(\xi_{\nu}))}{3\zeta(3)/2}\right),$
@@
The present neutrino abundance is further modified by spectral distortions of neutrinos and antineutrinos, though the integrals over their distributions cannot be carried out analytically in this case. The effect of primordial neutrino asymmetries becomes clearer when spectral distortions are neglected, allowing an analytical evaluation of the integrals, as shown in the second row of (96). The first factor, $\frac{\sum\limits_{i}m_{i}}{93.12\,{\rm eV}}$, represents the usual neutrino relic abundance in the SM, while the second factor accounts for the effects of neutrino asymmetries, with ${\rm Li}_{j}(x)$ being the polylogarithms of order $j$ and $\zeta(x)$ the Riemann zeta function. In the SM case, ${\rm Li}_{3}(-1)=-3\zeta(3)/4$, so the second factor equals 1. For non-zero $\xi_{\nu}$, the second factor depends on the absolute value of $\xi_{\nu}$, with larger $\xi_{\nu}$ leading to a larger factor and, consequently, a larger neutrino abundance. However, the corrections from primordial neutrino asymmetries are small. For instance, the relative difference in $\Omega_{\nu}h^{2}$ for $\xi_{\nu}=0.2$ is only $1.54\%$. Similarly, the corrections to the neutrino abundance from spectral distortions of neutrinos and antineutrinos are negligible.
@@ 5.1.2 Effects on the Neutrino Boltzmann Hierarchies
where $\tau\equiv\int dt/a(t)$ is the conformal time, $y\equiv p/T_{\rm CM}$ is the comoving momentum, $\hat{n}\equiv\vec{p}/p$ is the direction of the momentum, and $\varrho_{i}^{0}(y)$ represents the unperturbed zeroth-order neutrino density distributions obtained in Section 3.
@@
$\delta_{i}(\vec{x},t)\equiv\frac{\rho_{i}(\vec{x},t)-\rho_{i,0}(t)}{\rho_{i,0}(t)}=\frac{4\pi T_{\rm CM}^{4}\int dy~y^{2}\varrho_{i}^{0}(y)\Psi_{i,0}\epsilon}{\rho_{i,0}(t)},$ (105)
@@ 5.1.3 Effects on the BAO
$\dot{\delta}_{\gamma}$ $=-\frac{4}{3}\theta_{\gamma}-\frac{2}{3}\dot{h},$ (106)
@@
$\dot{F}_{\gamma\,2}$ $=2\dot{\sigma}_{\gamma}=\frac{8}{15}\theta_{\gamma}-\frac{3}{5}kF_{\gamma\,3}+\frac{4}{15}\dot{h}+\frac{8}{5}\dot{\eta}-\frac{9}{5}an_{e}\sigma_{T}\sigma_{\gamma}+\frac{1}{10}an_{e}\sigma_{T}(G_{\gamma\,0}+G_{\gamma\,2}),$
@@
$\dot{G}_{\gamma\,l}$ $=\frac{k}{2l+1}\left[lG_{\gamma\,(l-1)}-(l+1)G_{\gamma\,(l+1)}\right]+an_{e}\sigma_{T}\left[-G_{\gamma\,l}+\frac{1}{2}\left(F_{\gamma\,2}+G_{\gamma\,0}+G_{\gamma\,2}\right)\left(\delta_{l0}+\frac{\delta_{l2}}{5}\right)\right].$
@@ 6.1 Observational Constraints on Primordial Neutrino Asymmetry
[…]
@@
[…]
@@
[…]
@@
We find that constraints on other parameters, such as $\omega_{b}$, $\omega_{cdm}$, $n_{s}$, and $m_{\nu_{1}}$, also vary significantly depending on the dataset used, as shown in Fig. 12. For $\omega_{b}$, the discrepancy is mainly attributed to the choice of nuclear reaction rates in the PRIMAT code during the BBN process, which predicts a deuterium abundance lower than current constraints from PDG, leading to a significantly lower $\omega_{b}$ Pitrou:2018cgg ; Burns:2022hkq ; Escudero:2022okz . While constraints on the primordial neutrino asymmetry $\xi_{\nu}$ do not strongly depend on the choice of nuclear reaction rates, as discussed in Escudero:2022okz , a detailed study using different nuclear reaction rates will be presented in our companion paper Li:2025rjr . The inclusion of BOSS BAO data significantly modifies the constraints on $\omega_{cdm}$, $n_{s}$, and $m_{\nu_{1}}$ due to its sensitivity to the evolution of the late Universe. In particular, the $95\%$ upper limit on the total neutrino mass tightens from $0.414$ eV (BBN + CMB) to $0.131$ eV (BBN + CMB + BAO), as BAO data help break the CMB degeneracy. Recently, DESI BAO observations have placed even stronger constraints on neutrino masses, $\sum m_{\nu_{i}}\lesssim 0.07\,\mathrm{eV}$ DESI:2024mwx . Relevant studies on primordial neutrino asymmetry and neutrino masses using DESI BAO data will be addressed in our companion paper Li:2025rjr .
[…]
@@ 6.2 Constraints on UV Model Parameters
Following Kawasaki:2022hvx , the primordial lepton asymmetry $\eta_{L}$ generated by the decay of Q-balls can be written as [^4] The difference between the expression here and the expression in  Kawasaki:2022hvx is due to the different definition of the primordial lepton asymmetry $\eta_{L}$.
$\eta_{L}=\frac{3\pi^{4}g_{\star}T_{\mathrm{D}}}{495\zeta(3)m_{3/2}},$ (108)
where $g_{\star}=10.75$ is the relativistic degree of freedom at neutrino decoupling, $m_{3/2}$ is the gravitino mass and $T_{\mathrm{D}}$ is the L-balls’ decay temperature, which is a function of other UV model parameters:
@@
$\hskip 15.0pt\times\left(\frac{g_{*}}{10.75}\right)^{-1/4}\left(\frac{\beta}{6\times 10^{-4}}\right)^{-5/8}\left(\frac{m_{3/2}}{0.5~\mathrm{GeV}}\right)^{5/2}\left(\frac{M_{F}}{5\times 10^{6}~\mathrm{GeV}}\right)^{-2}\left(\frac{N_{\ell}}{3}\right)^{1/2}\left(\frac{\zeta}{2.5}\right)^{1/2},$ (109)
@@
$\simeq\left(\frac{m_{3/2}}{0.5~\mathrm{GeV}}\right)\,(1.65\pm 0.86)~\mathrm{MeV}\quad\,[\textbf{BBN+CMB+BAO}],$ (110)
@@
$\simeq\left(\frac{T_{\mathrm{D}}}{3.3~\mathrm{MeV}}\right)(1.00^{+1.08}_{-0.34})~\mathrm{GeV}\quad\,[\textbf{BBN+CMB+BAO}].$ (111)
@@ 7 Conclusion
In this work, we address these ambiguities by numerically solving the neutrino decoupling process and analyzing the resulting (anti)neutrino density matrices in the context of BBN, CMB, and LSS. We rederive the quantum kinetic equations (QKEs) for neutrinos and antineutrinos using the Closed Time Path (CTP) formalism for two-point functions. Focusing on the complete flavor equilibration cases with $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{\nu}$ , we numerically solve the QKEs in the early universe using an extended version of the public code FortEPiaNO Gariazzo:2019gyi ; Bennett:2020zkv . Our calculations include the effects of neutrino oscillations, matter interactions, FTQED corrections up to $\mathcal{O}(e^{3})$, and comprehensive neutrino-electron and neutrino-neutrino collision terms. The final $N_{\rm eff}$ is then computed based on the resulting density matrices and can be expressed as $N_{\rm eff}\simeq N_{\rm eff}^{\rm SM}+3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right)+0.0102\xi_{\nu}^{2}$, where $N_{\rm eff}^{\rm SM}=3.0440\pm 0.0002$. The second and third terms represent corrections from nonzero $\xi_{\nu}$ in the instantaneous decoupling limit and non-instantaneous effects, respectively. We also confirm that spectral distortions of neutrinos and antineutrinos with nonzero $\xi_{\nu}$ differ due to distinct Pauli-blocking effects, in agreement with preliminary studies Grohs:2016cuu . The formalism we develop can also be straightforwardly applied to the study of flavor equilibration processes, see Pastor:2008ti ; Mangano:2011ip ; Froustey:2024mgf for previous relevant discussions.
@@
[…]
@@ Acknowledgements.
We thank J. Froustey and C. Pitrou for helpful communications about their relevant work. This work is supported by the NSFC under Grants No. 12347105, No. 12375099 and No. 12447101, and the National Key Research and Development Program of China Grant No. 2020YFC2201501, No. 2021YFA0718304. The authors gratefully acknowledge the use of publicly available codes FortEPiaNO Gariazzo:2019gyi ; Bennett:2020zkv , PRIMAT Pitrou:2018cgg , CLASS Lesgourgues:2011re ; Lesgourgues:2011rh and MontePythonAudren:2012wb ; Brinckmann:2018cvx .
@@ A.1 Complete expressions for the collision terms
Finally, the collision terms for the antineutrinos $\bar{\mathcal{I}}\equiv\bar{\mathcal{I}}_{\rm sc}+\bar{\mathcal{I}}_{\rm ann}+\bar{\mathcal{I}}_{\nu\nu}$ can be obtained simply by the transformations $\varrho\leftrightarrow\bar{\varrho}$ and $L\leftrightarrow R$.
@@ A.2 Reduced expressions for the comoving collision terms
$\tilde{\mathcal{I}}_{\rm sc}$ $=$ $\int{\rm d}y_{2}{\rm d}y_{3}\frac{y_{2}}{E_{2}}$
@@
$\tilde{\mathcal{I}}_{\rm ann}$ $=$ $\int{\rm d}y_{2}{\rm d}y_{3}\frac{y_{3}}{E_{3}}$
@@
$\tilde{\mathcal{I}}_{\nu\nu}$ $=$ $\frac{1}{4}\int{\rm d}y_{2}{\rm d}y_{3}\,$
@@
Again, the comoving collision terms for the antineutrinos $\tilde{\bar{\mathcal{I}}}\equiv\tilde{\bar{\mathcal{I}}}_{\rm sc}+\tilde{\bar{\mathcal{I}}}_{\rm ann}+\tilde{\bar{\mathcal{I}}}_{\nu\nu}$ can be obtained simply by the transformations $\varrho\leftrightarrow\bar{\varrho}$ and $L\leftrightarrow R$.